OpenStudy (anonymous):

The graph above shows the function g, where g(x) =k(x+3)(x-3) for some constant k. If g(a-1.2)=0 and a is greater than 0, what is the value of a?

hartnn (hartnn):

g(3) and g(-3) are 0 so a-1.2 = 3 or a-1.2 = -3

OpenStudy (anonymous):



OpenStudy (anonymous):

4.2 or -1.8?

hartnn (hartnn):

since a>0 a=3+1.2

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so 4.2?